# Graph Kernel Neural Networks

### v2

@@ Abstract
Index Terms:
Graph neural network, graph kernel, deep learning
publicationid: pubid: $0$–$00$ 2021 IEEEpublicationid: pubid:
@@ I Introduction
In recent years, graph neural networks (GNNs) have gained increasing traction in the machine learning community. Graphs have long been used as a powerful abstraction for a wide variety of real-world data where structure plays a key role [1], from collaborations [2, 3] to biological [4, 5] and physical [6] data, to mention a few. Before the advent of GNNs, graph kernels provided a principled way to deal with graph data in the traditional machine learning setting [7, 8, 9]. However, with the attention of machine learning researchers steadily shifting away from hand-crafted features toward end-to-end models where both the features and the model are learned together, neural networks have quickly overtaken kernels as the framework of choice to deal with graph data, leading to multiple popular architectures such as [3, 4, 10].
@@
  • We introduce a neural model that is fully structural, unlike existing approaches that require embedding the input graph into a larger, relaxed space. While the latter allows one to seek more complex and arbitrary decision boundaries, it also has the potential of overfitting the problem and creating more local optima, as evidenced by the need to use dropout strategies to avoid overfitting in Graph Convolutional Networks (GCNs) [11, 12].
@@
  • Our architecture allows to plug-in any type of graph kernel (not just differentiable kernels as in [13, 14]).
@@ II Related work
The majority of graph kernels belong to one of two main categories: 1) bag-of-structures and 2) information propagation kernels. Bag-of-structures kernels compute the similarity between a pair of input graphs by first decomposing them into simpler substructures and then counting the number of isomorphic substructures between the two input graphs. Depending on the type of substructure considered, one can build a multitude of different kernels, e.g., subtrees [15], shortest paths [16], and graphlets [17]. Information propagation kernels, on the other hand, include methods where pairs of input graphs are compared based on how information diffuses on them. Examples include random walk kernels [18, 19], quantum walk kernels [8, 20, 21], and kernels based on iterative label refinements [7]. While some kernels work only on undirected and unattributed graphs, other kernels are designed to handle attributes as well, either discrete- or continuous-valued [7, 22]. For a detailed review and historical perspective on graph kernels, we refer the reader to the recent survey of Kriege et al.  [23].
@@
In recent years, with the advent of deep learning and the renewed interest in neural architectures, the focus of graph-based machine learning researchers has quickly moved to extending deep learning approaches to deal with graph data. Fundamentally, the principle underpinning most GNNs is that of exploiting the structure of the graph to propagate the node feature information iteratively. One of the first papers to propose the idea of GNNs is that of Scarselli et al.  [24], where an information diffusion mechanism is used to learn the nodes’ latent representations by exchanging neighborhood information. Sperduti and Starita [25] and Micheli [26] had previously relied instead on recursive operators, which can be seen as a variation of convolutional operators where different layers share the same weights, to adapt neural networks to operate on structured data such as graphs. Most modern GNNs, however, appear to fall within three main categories, as discussed by Bronstein et al.  [27]: 1) convolutional [3, 28, 29, 30, 31], 2) attentional [10], and 3) message passing [4]. The message-passing model is often seen as the most general one and has been shown to be formally equivalent to the WL graph isomorphism test under some technical conditions [32, 33].
@@
An increasing number of researchers are trying to overcome the latter limitation by moving away from standard message-passing GNNs [34, 35]. Eliasof et al.  [36] propose a reinterpretation of graph convolution in terms of partial differential equations on graphs, noting that different problems can benefit from different networks dynamics, hence suggesting the need to look beyond diffusion. Bevilaqua et al.  [37] represent graphs as bags of substructures and show that this allows the WL test to distinguish between otherwise indistinguishable graphs. Bouritsas et al.  [38] take instead an alternative approach where augmenting the graph nodes attributes with positional encoding is shown to improve the ability of the WL test to discriminate between non-isomorphic graphs. The field of GNNs is, in general, in rapid and continuous expansion, so presenting a comprehensive review of the literature is beyond the scope of this paper. Instead, we refer the interested readers to [39, 27].
@@
In this work, we argue that a natural extension of the convolution operation to the graph domain, and thus an ideal candidate to build graph CNNs, already exists in the form of graph kernels. A number of works in the literature have investigated potential synergies between GNNs and kernels. Lei et al.  [40] introduce a class of deep recurrent neural architectures to show that this lies in the reproducing kernel Hilbert space (RKHS) of graph kernels. Nikolentzos et al.  [41] compute continuous embeddings of graphs using kernels and plug them into a neural network. Xu et al.  [32] and Morris et al.  [33] show that GNNs have the same expressiveness as the WL graph kernel [7] and propose a new generalized architecture with increased expressive power. Du et al.  [42] take a somewhat different approach and exploit neural networks to introduce a new graph kernel. This, in turn, is shown to be equivalent to an infinitely wide GNN initialized with random weights and trained with gradient descent. Chen et al.  [43] propose a graph neural architecture where each layer enumerates local substructures around each node and then maps them to an RKHS via a Gaussian kernel mapping. In the context of graph compression, Bouritsas et al.  [44] learn how to best decompose the graph into small substructures that are also learned. Navarin et al.  [45] propose a modification of DGCNN [46], a popular GNN architecture, where a multi-task learning approach is used to drive the learned node embeddings to be close to those computed by graph kernels.
@@ III-A Preliminaries
Kernel methods. This class of algorithms is capable of learning when presented with a particular positive pairwise measure on the input data, known as a kernel. Consider a set $X$ and a positive semi-definite kernel function $\mathcal{K}:X\times X\rightarrow\mathbb{R}$ such that there exists a map $\phi:X\rightarrow H$ into a Hilbert space $H$ and $\mathcal{K}(x,y)=\phi(x)^{\top}\phi(y)$ for all $x,y\in X$. Crucially, $X$ can represent any set of data on which a kernel can be defined, from $\mathbb{R}^{d}$ to a finite set of graphs. Hence, the field of machine learning is ripe with examples of graph kernels (see Section II), which are nothing but positive semi-definite pairwise similarity measures [^1] Note that while several kernels can be interpreted as similarity measures, this is not always the case [47]. on graphs. These can be either implicit (only $\mathcal{K}$ is computed) or explicit ($\phi$ is also computed).
Weisfeiler-Lehman test. The WL kernel [7] is one of the most powerful and commonly used graph kernels, and it is based on the 1-dimensional Weisfeiler-Lehman (1-WL) graph isomorphism test. The idea underpinning the test is to partition the node set by iteratively propagating the node labels between adjacent nodes. With each iteration, the set of labels accumulated at each node of the graph is then mapped to a new label through a hash function. This procedure is repeated until the cardinality of the set of labels stops growing. Two graphs can then be compared in terms of their label sets at convergence, with two graphs being isomorphic only if their label sets coincide.
Learning on graphs. Let $\mathcal{G}=(V,X,E)$ be an undirected graph with $|V|$ nodes and $|E|$ edges, where each node $v$ is associated to a label $x(v)$ belonging to a dictionary $\mathbf{D}$. Note that, generally, graphs can be directed and have continuous and discrete attributes on both the nodes and edges; however, for simplicity in this paper, we restrict our attention to undirected node-labeled graphs. Indeed, the model described in this paper can be easily applied to directed, edge-labeled graphs as well. A common goal in graph machine learning problems is to produce a vector representation of $\mathcal{G}$ that is aware of both the node labels and the structural information of $\mathcal{G}$. Generalizing the convolution operator to graphs, GNNs learn the parameters $\Theta$ of a function $h$ that performs message passing [4] in the 1-hop neighborhood of each node $v\in V$,
@@ III-B Proposed architecture
Since in this paper we focus on graph-level tasks, the final graph embedding after $L$ layers is obtained by $\square(Z^{1}|\dots|Z^{L})$, where $Z^{\ell}$ are the output features of the $\ell$th layer and $\cdot$|$\cdot$ | ⋅ indicates the concatenation of node-wise features. This is then fed to a multilayer perceptron (MLP) layer to obtain the final classification. In particular, in all our experiments we use sum pooling as the aggregation operator $\square$. Figure 1 shows an overview of the proposed multilayer architecture.
@@
Note that while previous work in the literature used the JSD to generalize the widely used cross-entropy loss function [48], here we use it as a regularization term. That is, rather than minimizing the JSD or cross-entropy in order to limit the distance to a target class [49], we maximize the JSD to force the learned distributions to be as far away from one another as possible.
@@ III-D Expected kernel estimation through importance sampling
[…]
@@ III-E Expressive power
In [33], Morris et al. show that the expressive power of standard GNNs (i.e., GNNs that only consider the immediate neighbors of a node when updating the labels) is equivalent to that of the 1-WL test. In this subsection, we discuss our model’s expressive power and argue that this is higher than the 1-WL test and thus standard message-passing GNNs. For this to be the case, our model should be both able to distinguish every pair of graphs that can be distinguished by the 1-WL test and it should be able to distinguish between pairs of graphs where the 1-WL test fails.
Graphs distinguished by the 1-WL test. Consider two unlabelled graphs $\mathcal{G}_{1}=(V_{1},E_{1})$ and $\mathcal{G}_{2}=(V_{2},E_{2})$ with the same number of nodes $n=|V_{1}|=|V_{2}|$, such that the 1-WL test can distinguish between them. Then we can also build a GKNN able to distinguish between them by adopting the WL kernel in the GKC layer, and the radius $r$ is such that the subgraphs $N_{\mathcal{G}_{1}}^{r}(v)$ span the entire graph $\mathcal{G}_{1}$ (similarly for $\mathcal{G}_{2}$). In fact, the set of labels computed against any one mask will be the same for every node in graph $\mathcal{G}_{1}$ ($\mathcal{G}_{2}$) since all the subgraphs are identically equal to the graph itself. Assuming that we have a set of structural masks $\{\mathcal{M}_{1},\cdots,\mathcal{M}_{m}\}$ that span the space of graphs over $n$ nodes, then, under the hypothesis that the 1-WL test can distinguish between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, the set of labels computed by the GKC layer (i.e., the responses of the masks) while identical on all the nodes of each graph, will differ between the two graphs (see Figure 3). In particular, the response will peak for any mask fractionally equivalent to the graph itself [50]. Therefore, under any reasonable pooling strategy, the GKNN will be able to distinguish between the two graphs.
@@
Figure 4: $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ cannot be distinguished by the 1-WL test, however the GKNN correctly identifies them as non-isomorphic graphs. If $M$ is the set of all structural masks representing stars up to $n-1$ nodes, no star mask will achieve maximum similarity on the radius 1 subgraphs centered on the nodes of $\mathcal{G}_{1}$, while the same maximal response will be achieved when considering the radius 1 subgraphs centered on the nodes of $\mathcal{G}_{2}$.
@@
Graphs where the 1-WL test fails. Not only the GKNN can distinguish between pairs of graphs that are correctly distinguished by the 1-WL test, but it also succeeds where the 1-WL test fails. To see that this is the case, consider the standard example of two graphs with 6 nodes, where $\mathcal{G}_{1}$ is the dumbbell graph obtained by connecting two 3-nodes cycles with one edge, while $\mathcal{G}_{2}$ is a $2\times 3$ lattice graph (see Figure 4). Both graphs have the same degree distribution, with 4 nodes of degree 2 and 2 nodes of degree 3. As explained in [51, 52], despite being non-isomorphic, the two graphs are considered identical by the 1-WL test (in other words, the WL relabeling procedure converges to the same set of labels for the two graphs). To see why this happens, consider that through the propagation phases, each node is only made aware of what degree other nodes in the graph have, and not how many such nodes can be reached. However, we can easily see that this is not an issue for the GKNN.
@@
Assume we have a GKC layer with radius $r=1$ that adopts a WL kernel and a set of structural masks $M=\{\mathcal{M}_{1},\cdots,\mathcal{M}_{m}\}$ representing all stars up to $n-1$ nodes, where $n$ denotes the number of nodes of the input graph. The presence of a triangle in the graph induces the existence of at least three subgraphs of radius 1 with connections between neighbors, i.e., not stars. On these subgraphs, no star masks will achieve maximum similarity under the WL kernel. Having a pooling process that normalizes the masks responses with respect to size and maximizes over the masks on the same node, implies that each node will have the same maximal response if it does not contain triangles and a lower response otherwise. A min-pooling process over the node features will then be able to discriminate between graphs with and without triangles. Now consider the graphs $G_{1}$ and $G_{2}$ introduced above where the 1-WL test failed. As shown in Figure 4 the subgraphs built around the nodes of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ will be structurally different, i.e., triangles in $\mathcal{G}_{1}$ and 3 nodes path graphs in $\mathcal{G}_{2}$. As a consequence, the feature vectors $x(v)$ over the nodes of the two graphs will be different, allowing the GKNN to distinguish between them.

### v3

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I Introduction
In recent years, graph neural networks (GNNs) have gained increasing traction in the machine learning community. Graphs have long been used as a powerful abstraction for a wide variety of real-world data where structure plays a key role [duin2012dissimilarity], from collaborations [lima2014coding, kipf2016semi] to biological [gilmer2017neural, ye2015thermodynamic] and physical [shlomi2020graph] data, to mention a few. Before the advent of GNNs, graph kernels provided a principled way to deal with graph data in the traditional machine learning setting [shervashidze2011weisfeiler, bai2015quantum, minello2019can]. However, with the attention of machine learning researchers steadily shifting away from hand-crafted features toward end-to-end models where both the features and the model are learned together, neural networks have quickly overtaken kernels as the framework of choice to deal with graph data, leading to multiple popular architectures such as [kipf2016semi, gilmer2017neural, velickovic2018graph].
@@
  • We introduce a neural model that is fully structural, unlike existing approaches that require embedding the input graph into a larger, relaxed space. While the latter allows one to seek more complex and arbitrary decision boundaries, it also has the potential of overfitting the problem and creating more local optima, as evidenced by the need to use dropout strategies to avoid overfitting in Graph Convolutional Networks (GCNs) [rong2019dropedge, cong2021provable].
@@
  • Our architecture allows to plug-in any type of graph kernel (not just differentiable kernels as in [nikolentzos2020random, feng2022kergnns]).
@@ II Related work
The majority of graph kernels belong to one of two main categories: 1) bag-of-structures and 2) information propagation kernels. Bag-of-structures kernels compute the similarity between a pair of input graphs by first decomposing them into simpler substructures and then counting the number of isomorphic substructures between the two input graphs. Depending on the type of substructure considered, one can build a multitude of different kernels, e.g., subtrees [ramon2003expressivity], shortest paths [borgwardt2005shortest], and graphlets [shervashidze2009efficient]. Information propagation kernels, on the other hand, include methods where pairs of input graphs are compared based on how information diffuses on them. Examples include random walk kernels [kashima2003marginalized, bai2013graph], quantum walk kernels [bai2015quantum, rossi2015measuring, zhang2020r], and kernels based on iterative label refinements [shervashidze2011weisfeiler]. While some kernels work only on undirected and unattributed graphs, other kernels are designed to handle attributes as well, either discrete- or continuous-valued [shervashidze2011weisfeiler, da2017tree]. For a detailed review and historical perspective on graph kernels, we refer the reader to the recent survey of Kriege et al.  [kriege2020survey].
@@
In recent years, with the advent of deep learning and the renewed interest in neural architectures, the focus of graph-based machine learning researchers has quickly moved to extending deep learning approaches to deal with graph data. Fundamentally, the principle underpinning most GNNs is that of exploiting the structure of the graph to propagate the node feature information iteratively. One of the first papers to propose the idea of GNNs is that of Scarselli et al.  [scarselli2008graph], where an information diffusion mechanism is used to learn the nodes’ latent representations by exchanging neighborhood information. Sperduti and Starita [sperduti1997supervised] and Micheli [micheli2009neural] had previously relied instead on recursive operators, which can be seen as a variation of convolutional operators where different layers share the same weights, to adapt neural networks to operate on structured data such as graphs. Most modern GNNs, however, appear to fall within three main categories, as discussed by Bronstein et al.  [bronstein2021geometric]: 1) convolutional [kipf2016semi, atwood2016diffusion, levie2018cayleynets, bai2020learning, li2021learning], 2) attentional [velickovic2018graph], and 3) message passing [gilmer2017neural]. The message-passing model is often seen as the most general one and has been shown to be formally equivalent to the WL graph isomorphism test under some technical conditions [xu2018powerful, morris2019weisfeiler].
@@
An increasing number of researchers are trying to overcome the latter limitation by moving away from standard message-passing GNNs [bicciato2022classifying, bicciato2024gnn]. Eliasof et al.  [eliasof2021pde] propose a reinterpretation of graph convolution in terms of partial differential equations on graphs, noting that different problems can benefit from different networks dynamics, hence suggesting the need to look beyond diffusion. Bevilaqua et al.  [bevilacqua2022equivariant] represent graphs as bags of substructures and show that this allows the WL test to distinguish between otherwise indistinguishable graphs. Bouritsas et al.  [bouritsas2022improving] take instead an alternative approach where augmenting the graph nodes attributes with positional encoding is shown to improve the ability of the WL test to discriminate between non-isomorphic graphs. The field of GNNs is, in general, in rapid and continuous expansion, so presenting a comprehensive review of the literature is beyond the scope of this paper. Instead, we refer the interested readers to [wu2020comprehensive, bronstein2021geometric].
@@
In this work, we argue that a natural extension of the convolution operation to the graph domain, and thus an ideal candidate to build graph CNNs, already exists in the form of graph kernels. A number of works in the literature have investigated potential synergies between GNNs and kernels. Lei et al.  [lei2017deriving] introduce a class of deep recurrent neural architectures to show that this lies in the reproducing kernel Hilbert space (RKHS) of graph kernels. Nikolentzos et al.  [nikolentzos2018kernel] compute continuous embeddings of graphs using kernels and plug them into a neural network. Xu et al.  [xu2018powerful] and Morris et al.  [morris2019weisfeiler] show that GNNs have the same expressiveness as the WL graph kernel [shervashidze2011weisfeiler] and propose a new generalized architecture with increased expressive power. Du et al.  [du2019graph] take a somewhat different approach and exploit neural networks to introduce a new graph kernel. This, in turn, is shown to be equivalent to an infinitely wide GNN initialized with random weights and trained with gradient descent. Chen et al.  [chen2020convolutional] propose a graph neural architecture where each layer enumerates local substructures around each node and then maps them to an RKHS via a Gaussian kernel mapping. In the context of graph compression, Bouritsas et al.  [bouritsas2021partition] learn how to best decompose the graph into small substructures that are also learned. Navarin et al.  [navarin2020learning] propose a modification of DGCNN [zhang2018end], a popular GNN architecture, where a multi-task learning approach is used to drive the learned node embeddings to be close to those computed by graph kernels.
@@ III-A Preliminaries
Kernel methods. This class of algorithms is capable of learning when presented with a particular positive pairwise measure on the input data, known as a kernel. Consider a set $X$ and a positive semi-definite kernel function $\mathcal{K}:X\times X\rightarrow\mathbb{R}$ such that there exists a map $\phi:X\rightarrow H$ into a Hilbert space $H$ and $\mathcal{K}(x,y)=\phi(x)^{\top}\phi(y)$ for all $x,y\in X$. Crucially, $X$ can represent any set of data on which a kernel can be defined, from $\mathbb{R}^{d}$ to a finite set of graphs. Hence, the field of machine learning is ripe with examples of graph kernels (see Section II), which are nothing but positive semi-definite pairwise similarity measures [^1] Note that while several kernels can be interpreted as similarity measures, this is not always the case [nebel2017types]. on graphs. These can be either implicit (only $\mathcal{K}$ is computed) or explicit ($\phi$ is also computed).
Weisfeiler-Lehman test. The WL kernel [shervashidze2011weisfeiler] is one of the most powerful and commonly used graph kernels, and it is based on the 1-dimensional Weisfeiler-Lehman (1-WL) graph isomorphism test. The idea underpinning the test is to partition the node set by iteratively propagating the node labels between adjacent nodes. With each iteration, the set of labels accumulated at each node of the graph is then mapped to a new label through a hash function. This procedure is repeated until the cardinality of the set of labels stops growing. Two graphs can then be compared in terms of their label sets at convergence, with two graphs being isomorphic only if their label sets coincide.
Learning on graphs. Let $\mathcal{G}=(V,X,E)$ be an undirected graph with $|V|$ nodes and $|E|$ edges, where each node $v$ is associated to a label $x(v)$ belonging to a dictionary $\mathbf{D}$. Note that, generally, graphs can be directed and have continuous and discrete attributes on both the nodes and edges; however, for simplicity in this paper, we restrict our attention to undirected node-labeled graphs. Indeed, the model described in this paper can be easily applied to directed, edge-labeled graphs as well. A common goal in graph machine learning problems is to produce a vector representation of $\mathcal{G}$ that is aware of both the node labels and the structural information of $\mathcal{G}$. Generalizing the convolution operator to graphs, GNNs learn the parameters $\Theta$ of a function $h$ that performs message passing [gilmer2017neural] in the 1-hop neighborhood of each node $v\in V$,
@@ III-B Proposed architecture
[…]
@@
Note that while previous work in the literature used the JSD to generalize the widely used cross-entropy loss function [englesson2021generalized], here we use it as a regularization term. That is, rather than minimizing the JSD or cross-entropy in order to limit the distance to a target class [bai2019quantum], we maximize the JSD to force the learned distributions to be as far away from one another as possible.
@@ III-D Expected kernel estimation through importance sampling
$\mathbb{E}_{p}[f]=\int_{X}f(x)\,p(x)dx=\int_{X}f(x)\frac{p(x)}{g(x)}\,g(x)dx=\\ \mathbb{E}_{g}\left[f(x)\frac{p(x)}{g(x)}\right]\,.$ (8)
@@ III-E Expressive power
In [morris2019weisfeiler], Morris et al. show that the expressive power of standard GNNs (i.e., GNNs that only consider the immediate neighbors of a node when updating the labels) is equivalent to that of the 1-WL test. In this subsection, we discuss our model’s expressive power and argue that this is higher than the 1-WL test and thus standard message-passing GNNs. For this to be the case, our model should be both able to distinguish every pair of graphs that can be distinguished by the 1-WL test and it should be able to distinguish between pairs of graphs where the 1-WL test fails.
Graphs distinguished by the 1-WL test. Consider two unlabelled graphs $\mathcal{G}_{1}=(V_{1},E_{1})$ and $\mathcal{G}_{2}=(V_{2},E_{2})$ with the same number of nodes $n=|V_{1}|=|V_{2}|$, such that the 1-WL test can distinguish between them. Then we can also build a GKNN able to distinguish between them by adopting the WL kernel in the GKC layer, and the radius $r$ is such that the subgraphs $N_{\mathcal{G}_{1}}^{r}(v)$ span the entire graph $\mathcal{G}_{1}$ (similarly for $\mathcal{G}_{2}$). In fact, the set of labels computed against any one mask will be the same for every node in graph $\mathcal{G}_{1}$ ($\mathcal{G}_{2}$) since all the subgraphs are identically equal to the graph itself. Assuming that we have a set of structural masks $\{\mathcal{M}_{1},\cdots,\mathcal{M}_{m}\}$ that span the space of graphs over $n$ nodes, then, under the hypothesis that the 1-WL test can distinguish between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, the set of labels computed by the GKC layer (i.e., the responses of the masks) while identical on all the nodes of each graph, will differ between the two graphs (see Figure 3). In particular, the response will peak for any mask fractionally equivalent to the graph itself [ramana1994fractional]. Therefore, under any reasonable pooling strategy, the GKNN will be able to distinguish between the two graphs.
@@
Figure 4: $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ cannot be distinguished by the 1-WL test, however the GKNN correctly identifies them as non-isomorphic graphs. If $M$ is the set of all structural masks representing stars up to n1n-1 nodes, no star mask will achieve maximum similarity on the radius 1 subgraphs centered on the nodes of $\mathcal{G}_{1}$, while the same maximal response will be achieved when considering the radius 1 subgraphs centered on the nodes of $\mathcal{G}_{2}$.
@@
Graphs where the 1-WL test fails. Not only the GKNN can distinguish between pairs of graphs that are correctly distinguished by the 1-WL test, but it also succeeds where the 1-WL test fails. To see that this is the case, consider the standard example of two graphs with 6 nodes, where $\mathcal{G}_{1}$ is the dumbbell graph obtained by connecting two 3-nodes cycles with one edge, while $\mathcal{G}_{2}$ is a $2\times 3$ lattice graph (see Figure 4). Both graphs have the same degree distribution, with 4 nodes of degree 2 and 2 nodes of degree 3. As explained in [sato2020survey, nikolentzos2020k], despite being non-isomorphic, the two graphs are considered identical by the 1-WL test (in other words, the WL relabeling procedure converges to the same set of labels for the two graphs). To see why this happens, consider that through the propagation phases, each node is only made aware of what degree other nodes in the graph have, and not how many such nodes can be reached. However, we can easily see that this is not an issue for the GKNN.
@@
Assume we have a GKC layer with radius $r=1$ that adopts a WL kernel and a set of structural masks $M=\{\mathcal{M}_{1},\cdots,\mathcal{M}_{m}\}$ representing all stars up to n1n-1 nodes, where $n$ denotes the number of nodes of the input graph. The presence of a triangle in the graph induces the existence of at least three subgraphs of radius 1 with connections between neighbors, i.e., not stars. On these subgraphs, no star masks will achieve maximum similarity under the WL kernel. Having a pooling process that normalizes the masks responses with respect to size and maximizes over the masks on the same node, implies that each node will have the same maximal response if it does not contain triangles and a lower response otherwise. A min-pooling process over the node features will then be able to discriminate between graphs with and without triangles. Now consider the graphs $G_{1}$ and $G_{2}$ introduced above where the 1-WL test failed. As shown in Figure 4 the subgraphs built around the nodes of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ will be structurally different, i.e., triangles in $\mathcal{G}_{1}$ and 3 nodes path graphs in $\mathcal{G}_{2}$. As a consequence, the feature vectors $x(v)$ over the nodes of the two graphs will be different, allowing the GKNN to distinguish between them.